# The TeraZ Mirage: New Physics Lost in Blind Directions

### v1

@@ Abstract
The next generation of high-luminosity electron-positron colliders, such as FCC-ee and CEPC operating at the $Z$ pole (TeraZ), is expected to deliver unprecedented precision in electroweak measurements. These precision observables are typically interpreted within the Standard Model Effective Field Theory (SMEFT), offering a powerful tool to constrain new physics. However, the large number of independent SMEFT operators allows for the possibility of blind directions, parameter combinations to which electroweak precision data are largely insensitive. In this work, we demonstrate that such blind directions are not merely an artefact of agnostic effective field theory scans, but arise generically in realistic ultraviolet completions involving multiple heavy fields. We identify several concrete multi-field extensions of the Standard Model whose low-energy SMEFT projections align with known blind subspaces, and show that these persist even after accounting for renormalisation group evolution and finite one-loop matching effects. Our analysis highlights that the apparent sensitivity to new physics of TeraZ may be significantly overestimated, and that indirect searches alone are often insufficient to rule out broad classes of ultraviolet physics. Complementary high-energy collider probes are therefore essential to fully explore the SMEFT parameter space.
@@ I Introduction
The next generation of circular electron-positron colliders, such as the proposed Future Circular Collider (FCC-ee) and the Circular Electron-Positron Collider (CEPC), have the capability to operate at the $Z$-boson threshold, producing an unprecedented number of $Z$ bosons. This high-statistics environment enables precise measurements of electroweak precision observables (EWPOs), allowing the test of the Standard Model (SM) with high sensitivity. Due to the high luminosity of these experiments, TeraZ, referring to the production of approximately $10^{9}$ $Z$ bosons, offers a window into potential deviations from the SM, thereby serving as a crucial indirect probe of new physics.
@@
Recent studies have highlighted the power of EWPO measurements in setting stringent and model-independent constraints on physics beyond the SM (BSM) [1, 2, 3]. It is argued that the precision nature of these observables makes them an essential tool in the search for new physics, as any deviation from SM predictions could signal the presence of new interactions or particles. Historically, a scientific outcome of the Large Electron-Positron Collider (LEP) was that such measurements provide sensitive probes to some BSM physics scenarios, guiding theoretical developments and placing constraints on new physics models [^1] A well-known counterexample to the claimed sensitivity of LEP-measured EWPOs to new physics is the scenario of a sequential fourth generation of fermions. In this case, the presence of fourth-generation fermions with masses of a few hundred GeV —i.e., not far above the electron-positron collision energy— could not be excluded due to a blind (or insensitive) direction in the $S-T$ parameter plane [4]. Thus, only Higgs boson measurements excluded a fourth generation at the LHC..
@@
Nowadays, the standard approach to interpreting small deviations from the SM is the framework of Effective Field Theory (EFT), which provides a systematic method for parametrising BSM effects in a model-independent manner. The power of indirect searches has been well established in flavour physics, where the Low-Energy Effective Field Theory (LEFT) framework has been instrumental in placing strong constraints on BSM scenarios [5, 6, 7]. However, when moving to the SMEFT, the complexity of the problem increases significantly. Unlike LEFT, which operates at lower energies with a relatively small operator basis, SMEFT introduces a vastly larger operator basis, comprising 2,499 independent dimension-six operators [8], many of which contribute to the 26 measured EWPOs. Moreover, this high-dimensional parameter space introduces the possibility of cancellations, where multiple SMEFT operators can conspire to mask new physics contributions, leading to potential blind spots in indirect searches. [^2] High-energy hadron colliders possess the advantage to be able to break such blind directions by exploiting different kinematic regions of the phase space, where the operators with non-trivial energy dependence contribute over proportionally compared to the SM, therefore being constrained more tightly [9, 10, 11, 12, 13].
@@
Addressing this question requires a deeper understanding of how concrete ultraviolet (UV) models project onto the SMEFT parameter space. To our knowledge, previous work has not addressed this question comprehensively. While it was acknowledged that a completely agnostic variation of EFT operator contributions to EWPOs leads to a plethora of blind directions [2], how pronounced they are in more realistic UV models has received little attention. Often, only a single field was integrated out, resulting in a small set of highly correlated EFT operators contributing to the EWPOs in the absence of blind directions [14, 2, 1, 3]. Consequently, testing only such scenarios overestimates the prowess of the TeraZ programme on new physics, as realistic UV models generically require the introduction of many new degrees of freedom. In particular, multi-field extensions of the SM may exhibit specific correlations among SMEFT coefficients that could lead to a “nightmare scenario” where new physics resides at energy scales not far away from HL-LHC energy ranges yet invisible to TeraZ. In this work, we make a first investigation of these correlations by exploring simple extensions of the SM and assessing whether they can evade detection at TeraZ. By doing so, we aim to quantify the extent to which TeraZ measurements can generically constrain new physics and to evaluate whether indirect detection remains a robust avenue for probing BSM physics.
A fully comprehensive and exhaustive treatment of all technical aspects is beyond the scope of this work. We introduce controlled simplifications in specific technical steps to maintain clarity in our argument. However, it is crucial to emphasise that relaxing any of these assumptions would only exacerbate the problem of blind directions in EWPO constraints. First, in our analysis, we refrain from marginalising over all effective operators. While such an approach would further expand the region of blind directions, it could also obscure the underlying correlations that relate these blind spots to the dynamics of UV model realisations. Second, we restrict to the sub-sector of the SMEFT spanned by four-fermion interactions with only third-family couplings, and we avoid blind directions that involve unreasonably different sizes of Wilson coefficients (WCs).
In Section II, we display the resulting blind directions of the so-called bottom-up approach, i.e. agnostically varying EFT operators contributing to the EWPOs. Then, in Section III, we consider which UV completions of the SMEFT generate the different blind directions upon integrating the new heavy fields out at tree level. In Section IV, we discuss the fit of selected scenarios to EWPO, including one-loop matching corrections. We offer a summary and conclusions in Section V.
@@ II Effective Field Theory Fit: bottom up
In this section, we adopt a bottom-up EFT approach to explore potential blind directions in EWPOs, assuming minimal prior knowledge about the UV origin of the effective operators. We also assess the impact of renormalisation group evolution (RGE), incorporating leading-log RGEs and full numerical integration of the one-loop SMEFT RGEs [15, 16, 17].
@@
[…]
@@
which represent the total decay width of the $W$, the partial width of the $W$ into different leptons, the partial width of the $W$ into quarks, the total $Z$ hadronic cross section, the $Z$ total decay rate, the forward-backward and left-right asymmetries as well as the ratios of the $Z$ decays into different quarks and leptons and the fine-structure constant. The definitions of these parameters and experimental values can be found in Refs. [18, 19, 20, 21, 22, 23, 24]. We take the Fermi constant $G_{F}$ and the $W$ mass $m_{W}$ as input parameters.
@@
When restricting to CP-conserving, baryon-number-preserving operators in the Warsaw basis [8] and neglecting flavour structure, we find $\mathcal{O}(50)$ operators, of which 10 independent operators can contribute to EWPOs at tree level [25]. Including RGE evolution, the number of operators contributing to EWPOs grows to 31, out of which 25 can arise from integrating out heavy fields at tree level in realistic UV completions of the SM. While it seems that the 26 EWPOs could overconstrain such a system, not all operators contribute to all EWPOs, and, thus, even imposing such strong constraints on the operators, one finds rather unconstrained subspaces of operators that lead to blind directions. For the numerical results of the fits, we use current constraints on the EWPOs, as implemented in smelli [26, 27], rather than relying on any TeraZ projection for their final uncertainties. While such a projection can have a significant effect on the size of the unconstrained region, it will not affect the existence of a blind direction.
@@
We focus on the space spanned by four-fermion operators, [^3] Note that blind directions within the rest of the SMEFT include the obvious cases of $\mathcal{O}_{\phi}$ —which affects only the Higgs potential [28]—, $\mathcal{O}_{\phi\square}$ —which enters weakly via RGEs— as well as certain operators that arise only at loop level. Note also that the combination $c_{\phi q}^{(1)}-c_{\phi q}^{(3)}$ is untestable with EWPO but only if it holds right at the $Z$ pole; RGE removes this one. as they are ubiquitous in models of new physics, arise easily at tree level in UV completions of the SMEFT and, with the exception of $\mathcal{O}_{ll}$ with first family leptons, none enters into EWPO at tree level [25]. The most unconstrained four-fermions involve third-family fermions only, in which we concentrate. Theoretically, one can motivate that new physics couples mostly to the third generation [29, 30, 31, 32]. We also ignore $c_{quqd}^{(1)}$, $c_{quqd}^{(8)}$, $c_{ledq}$, $c_{lequ}^{(1)}$ and $c_{lequ}^{(3)}$, as they are in general flavour-violating. We follow the notation of Ref. [8] for SMEFT WCs.
@@
We assume WCs to be defined at a scale $\Lambda=1$ TeV, and study their impact on EWPO at the scale $\sim m_{Z}$ using the leading-log dimension-six SMEFT RGEs, as well as upon integrating numerically these RGEs. To this latter goal, we rely on smelli which is further based on flavio [33] and Wilson [34].
@@
To start with, the following WCs do not contribute to EWPO or their contribution is negligible: $c_{uu}$, $c_{dd}$, $c_{qu}^{(8)}$, $c_{qd}^{(8)}$, $c_{ud}^{(8)}$, $c_{le}$, $c_{ll}$, $c_{ed}$, $c_{ld}$. This explains why most of these WCs do not appear in Fig. 1 of Ref. [2]. Out of these, though, $c_{uu}$ is known to be well constrained by two-loop effects captured by RGE re-summation [35, 36, 37, 2]; see Fig. 1.
@@
where we have included the most constraining EWPOs on these WCs. The null vector of the corresponding matrix, which describes the blind direction in this sub-space, is given by: $c_{ud}^{(1)}\sim c_{qd}^{(1)}$. It extends to the full set of EWPOs.
@@
$\Xi_{1}$ $(1,3)_{1}$ $+\frac{|y_{\Xi_{1}}|^{2}}{M_{\Xi_{1}}^{2}}\mathcal{O}_{ll}$
$\varphi$ $(1,2)_{1/2}$ $-\frac{|y^{e}_{\varphi}|^{2}}{2M_{\varphi}^{2}}\mathcal{O}_{le}-\frac{|y^{d}_{\varphi}|^{2}}{6M_{\varphi}^{2}}\left(\mathcal{O}^{(1)}_{qd}+6\mathcal{O}^{(8)}_{qd}\right)$
$\omega_{1}$ $(3,1)_{-1/3}$ $+\frac{|y^{du}_{\omega_{1}}|^{2}}{3M_{\omega_{1}}^{2}}\left(\mathcal{O}^{(1)}_{ud}-3\mathcal{O}^{(8)}_{ud}\right)+\frac{|y^{eu}_{\omega_{1}}|^{2}}{2M_{\omega_{1}}^{2}}\mathcal{O}_{eu}$
$+\frac{|y^{ql}_{\omega_{1}}|^{2}}{4M_{\omega_{1}}^{2}}\left(\mathcal{O}^{(1)}_{lq}-\mathcal{O}^{(3)}_{lq}\right)$
$\omega_{2}$ $(3,1)_{2/3}$ $+\frac{|y_{\omega_{2}}|^{2}}{M_{\omega_{2}}^{2}}\mathcal{O}_{dd}$
$\omega_{4}$ $(3,1)_{-4/3}$ $+\frac{|y^{ed}_{\omega_{4}}|^{2}}{2M_{\omega_{4}}^{2}\mathcal{O}_{ed}}$
$\Pi_{1}$ $(3,2)_{1/6}$ $-\frac{|y_{\Pi_{1}}|^{2}}{2M_{\Pi_{1}}^{2}}\mathcal{O}_{ld}$
$\Pi_{7}$ $(3,2)_{7/6}$ $-\frac{|y^{eq}_{\Pi_{7}}|^{2}}{2M_{\Pi_{7}}^{2}}\mathcal{O}_{qe}-\frac{|y^{lu}_{\Pi_{7}}|^{2}}{2M_{\Pi_{7}}^{2}}\mathcal{O}_{lu}$
$\zeta$ $(3,3)_{-1/3}$ $+\frac{|y^{ql}_{\zeta}|^{2}}{4M_{\zeta}^{2}}\left(3\mathcal{O}^{(1)}_{lq}+\mathcal{O}^{(3)}_{lq}\right)$
$\Omega_{1}$ $(6,1)_{1/3}$ $+\frac{|y^{ud}_{\Omega_{1}}|^{2}}{6M_{\Omega_{1}}^{2}}\left(2\mathcal{O}^{(1)}_{ud}+3\mathcal{O}^{(8)}_{ud}\right)$
$\Omega_{2}$ $(6,1)_{-2/3}$ $+\frac{|y_{\Omega_{2}}|^{2}}{2M_{\Omega_{2}}^{2}}\mathcal{O}_{dd}$
$\Phi$ $(8,2)_{1/2}$ $-\frac{|y^{dq}_{\Phi}|^{2}}{18M_{\Phi}^{2}}\left(4\mathcal{O}^{(1)}_{qd}-3\mathcal{O}^{(8)}_{qd}\right)$
@@
$\mathcal{B}$ $(1,1)_{0}$ $-\frac{|g^{d}_{\mathcal{B}}|^{2}}{2M_{\mathcal{B}}^{2}}\mathcal{O}_{dd}\,-\frac{|g^{l}_{\mathcal{B}}|^{2}}{2M_{\mathcal{B}}^{2}}\mathcal{O}_{ll}$
$\mathcal{B}_{1}$ $(1,1)_{1}$ $-\frac{|g^{du}_{\mathcal{B}_{1}}|^{2}}{3M_{\mathcal{B}_{1}}^{2}}\left(\mathcal{O}^{(1)}_{ud}+6\mathcal{O}^{(8)}_{ud}\right)$
$\mathcal{W}$ $(1,3)_{0}$ $-\frac{|g^{l}_{\mathcal{W}}|^{2}}{8M_{\mathcal{W}}^{2}}\mathcal{O}_{ll}$
$\mathcal{G}$ $(8,1)_{0}$ $-\frac{|g^{d}_{\mathcal{G}}|^{2}}{4M_{\mathcal{G}}^{2}}\mathcal{O}_{dd}$
$\mathcal{G}_{1}$ $(8,1)_{1}$ $-\frac{|g_{\mathcal{G}_{1}}|^{2}}{9M_{\mathcal{G}_{1}}^{2}}\left(4\mathcal{O}^{(1)}_{ud}-3\mathcal{O}^{(8)}_{ud}\right)$
$\mathcal{L}_{3}$ $(1,2)_{-3/2}$ $+\frac{|g_{\mathcal{L}_{3}}|^{2}}{M_{\mathcal{L}_{3}}^{2}}\mathcal{O}_{le}$
$\mathcal{U}_{2}$ $(3,1)_{2/3}$ $-\frac{|g^{ed}|^{2}}{M_{\mathcal{U}_{2}}^{2}}\mathcal{O}_{ed}-\frac{|g^{lq}_{\mathcal{U}_{2}}|^{2}}{2M_{\mathcal{U}_{2}}^{2}}\left(\mathcal{O}^{(1)}_{lq}+\mathcal{O}^{(3)}_{lq}\right)$
$\mathcal{U}_{5}$ $(3,1)_{5/3}$ $-\frac{|g_{\mathcal{U}_{5}}|^{2}}{M_{\mathcal{U}_{5}}^{2}}\mathcal{O}_{eu}$
$\mathcal{Q}_{1}$ $(3,2)_{1/6}$ $+\frac{2|g^{dq}_{\mathcal{Q}_{1}}|^{2}}{3M_{\mathcal{Q}_{1}}^{2}}\left(\mathcal{O}^{(1)}_{qd}-3\mathcal{O}^{(8)}_{qd}\right)+\frac{|g^{ul}_{\mathcal{Q}_{1}}|^{2}}{M_{\mathcal{Q}_{1}}^{2}}\mathcal{O}_{lu}$
$\mathcal{Q}_{5}$ $(3,2)_{-5/6}$ $+\frac{|g^{dl}_{\mathcal{Q}_{5}}|^{2}}{M_{\mathcal{Q}_{5}}^{2}}\mathcal{O}_{ld}+\frac{|g^{eq}_{\mathcal{Q}_{5}}|^{2}}{M_{\mathcal{Q}_{5}}^{2}}\mathcal{O}_{qe}$
$\mathcal{X}$ $(3,3)_{2/3}$ $-\frac{|g_{\mathcal{X}}|^{2}}{8M_{\mathcal{X}}^{2}}\left(3\mathcal{O}^{(1)}_{lq}-\mathcal{O}^{(3)}_{lq}\right)$
$\mathcal{Y}_{1}$ $(\bar{6},2)_{1/6}$ $+\frac{|g_{\mathcal{Y}_{1}}|^{2}}{3M_{\mathcal{Y}_{5}}^{2}}\left(2\mathcal{O}^{(1)}_{qd}+3\mathcal{O}^{(8)}_{qd}\right)$
@@
[…]

### v2

# Potential Blind Directions at TeraZ

@@ Abstract
The next generation of high-luminosity electron-positron colliders, such as FCC-ee and CEPC operating at the $Z$ pole (TeraZ), is expected to deliver unprecedented precision in electroweak measurements. These precision observables are typically interpreted within the Standard Model Effective Field Theory (SMEFT), offering a powerful tool to constrain new physics. However, the large number of independent SMEFT operators allows for the possibility of blind directions, parameter combinations to which electroweak precision data are largely insensitive. In this work, we demonstrate that such blind directions are not merely an artefact of agnostic effective field theory scans, but arise generically in realistic ultraviolet completions involving multiple heavy fields. We identify several concrete multi-field extensions of the Standard Model whose low-energy SMEFT projections align with known blind subspaces, and show that these persist even after accounting for renormalisation group evolution and finite one-loop matching effects. Our analysis shows that TeraZ will set a new benchmark in precision for indirect searches, but fully probing the space of possible ultraviolet physics requires moving towards a high energy hadron collider. Later FCC-ee runs at higher centre-of-mass energies, together with the FCC-hh, will provide the necessary complementary probes, enabling a far more complete exploration of the SMEFT parameter space.
@@ I Introduction
The next generation of circular electron-positron colliders, such as the proposed Future Circular Collider (FCC-ee) and the Circular Electron-Positron Collider (CEPC), have the capability to operate at the $Z$-boson threshold, producing an unprecedented number of $Z$ bosons. This high-statistics environment enables precise measurements of electroweak precision observables (EWPOs), allowing for the high-sensitivity testing of the Standard Model (SM). Due to the high luminosity of these experiments, TeraZ, referring to the production of approximately $10^{12}$ $Z$-bosons, offers a window into potential deviations from the SM, thereby serving as a crucial indirect probe of new physics.
@@
Recent studies have highlighted the power of EWPO measurements in setting stringent and model-independent constraints on physics beyond the SM (BSM) [Gargalionis:2024jaw, Maura:2024zxz, Allwicher:2024sso]. It is argued that the precision nature of these observables makes them an essential tool in the search for new physics, as any deviation from SM predictions could signal the presence of new interactions or particles. Historically, a scientific outcome of the Large Electron-Positron Collider (LEP) was that such measurements provide sensitive probes to some BSM physics scenarios, guiding theoretical developments and placing constraints on new physics models [^1] A well-known counterexample to the claimed sensitivity of LEP-measured EWPOs to new physics is the scenario of a sequential fourth generation of fermions. In this case, the presence of fourth-generation fermions with masses of a few hundred GeV —i.e., not far above the electron-positron collision energy— could not be excluded due to a blind (or insensitive) direction in the STS-T parameter plane [Kribs:2007nz]. Thus, only Higgs boson measurements excluded a fourth generation at the LHC..
@@
Nowadays, the standard approach to interpreting small deviations from the SM is the framework of Effective Field Theory (EFT), which provides a systematic method for parametrising BSM effects in a model-independent manner. The power of indirect searches has been well established in flavour physics, where the Low-Energy Effective Field Theory (LEFT) framework has been instrumental in placing strong constraints on BSM scenarios [Buchalla:1995vs, Buras:2000dm, Buttazzo:2017ixm]. However, when moving to the SMEFT, the complexity of the problem increases significantly. Unlike LEFT, which operates at lower energies with a relatively small operator basis, SMEFT introduces a vastly larger operator basis, comprising 2,499 independent dimension-six operators [Grzadkowski:2010es], many of which contribute to the 26 measured EWPOs. Moreover, this high-dimensional parameter space introduces the possibility of cancellations, where multiple SMEFT operators can conspire to mask new physics contributions, leading to potential blind spots in indirect searches. [^2] High-energy hadron colliders possess the advantage to be able to break such blind directions by exploiting different kinematic regions of the phase space, where the operators with non-trivial energy dependence contribute overproportionally compared to the SM, therefore being constrained more tightly [Englert:2015hrx, Englert:2017aqb, Franceschini:2017xkh, Banerjee:2018bio, Belvedere:2024wzg, Davighi:2024syj].
@@
Addressing this question requires a deeper understanding of how concrete ultraviolet (UV) models project onto the SMEFT parameter space. To the best of our knowledge, this question has not yet been explored in a fully comprehensive manner. While it has been recognized that a fully agnostic variation of EFT operator contributions to EWPOs gives rise to blind directions [Maura:2024zxz], their relevance in more realistic UV scenarios has received comparatively little attention. In many cases, only a single field has been integrated out, leading to a restricted set of tightly correlated EFT operators contributing to the EWPOs, and thus avoiding blind directions. [DasBakshi:2020pbf, Maura:2024zxz, Gargalionis:2024jaw, Allwicher:2024sso]. Consequently, considering only such simplified scenarios may not fully capture the reach of the TeraZ programme, since realistic UV models typically involve several new degrees of freedom. In particular, multi-field extensions of the SM can generate correlations among SMEFT coefficients that give rise to situations where new physics lies at scales close to those probed by the HL-LHC, yet remains difficult to access at TeraZ. In this work, we make a first investigation of these correlations by exploring simple extensions of the SM and assessing whether they can evade detection at TeraZ. By doing so, we aim to quantify the extent to which TeraZ measurements can generically constrain new physics and to evaluate whether indirect detection remains a robust avenue for probing BSM physics.
A fully comprehensive and exhaustive treatment of all technical aspects is beyond the scope of this work. We introduce controlled simplifications in specific technical steps to maintain clarity in our argument. It is worth noting, though, that relaxing any of these assumptions would generally worsen the issue of blind directions in EWPO constraints. First, in our analysis, we refrain from marginalising over all effective operators since otherwise it could obscure the underlying correlations that relate blind spots to the dynamics of UV model realisations. Second, we restrict to the sub-sector of the SMEFT spanned by four-fermion interactions with only third-family couplings, and we avoid exploring blind directions that involve unreasonably different sizes of Wilson coefficients (WCs).
Starting with LEP data first, in Section II we display the resulting blind directions of the so-called bottom-up approach, i.e. agnostically varying EFT operators contributing to the EWPOs. Then, in Section III, we consider which UV completions of the SMEFT generate the different blind directions upon integrating the new heavy fields out at tree level. In Section IV, we discuss the fit of selected scenarios to EWPO, including one-loop matching corrections. In Section V, we examine how these considerations evolve at TeraZ. We offer a summary and conclusions in Section VI.
@@ II Effective Field Theory Fit: bottom up
In this section, we adopt a bottom-up EFT approach to explore potential blind directions in EWPOs from LEP, assuming minimal prior knowledge about the UV origin of the effective operators. We also assess the impact of renormalisation group evolution (RGE), incorporating leading-log RGEs and full numerical integration of the one-loop SMEFT RGEs [Jenkins:2013wua, Jenkins:2013zja, Alonso:2013hga].
@@
$\text{EWPO}=\bigg\{$ $\Gamma_{W},\Gamma_{W}^{e\nu,\mu\nu,\tau\nu},\Gamma^{\text{had}}_{W},\sigma_{\text{had}},\Gamma_{Z},A_{\text{FB}}^{e,\mu,\tau},$
@@
which represent the total decay width of the $W$, the partial width of the $W$ into different leptons, the partial width of the $W$ into quarks, the total $Z$ hadronic cross section, the $Z$ total decay rate, the forward-backward and left-right asymmetries as well as the ratios of the $Z$ decays into different quarks and leptons and the fine-structure constant. The definitions of these parameters and experimental values can be found in Refs. [ATLAS:2014wva, Davier:2010nc, Group:2012gb, ALEPH:2005ab, ALEPH:2010aa, ParticleDataGroup:2024cfk, Biekotter:2025nln]. We take the Fermi constant $G_{F}$ and the $W$ mass $m_{W}$ as input parameters.
@@
When restricting to CP-conserving, baryon-number-preserving operators in the Warsaw basis [Grzadkowski:2010es] and neglecting flavour structure, we find $\mathcal{O}(50)$ operators, of which 10 independent operators can contribute to EWPOs at tree level [deBlas:2015aea]. Including RGE evolution, the number of operators contributing to EWPOs grows to 31, out of which 25 can arise from integrating out heavy fields at tree level in realistic UV completions of the SM. While it seems that the 26 EWPOs could overconstrain such a system, not all operators contribute to all EWPOs, and, thus, even imposing such strong constraints on the operators, one finds rather unconstrained subspaces of operators that lead to blind directions. For the numerical results of the fits, we use current constraints on the EWPOs, as implemented in smelli [Aebischer:2018iyb, Stangl:2020lbh].
@@
We focus on the space spanned by four-fermion operators, [^3] Note that blind directions within the rest of the SMEFT include the obvious cases of $\mathcal{O}_{\phi}$ —which affects only the Higgs potential [deBlas:2014mba]—, $\mathcal{O}_{\phi\square}$ —which enters weakly via RGEs— as well as certain operators that arise only at loop level. Note also that the combination $c_{\phi q}^{(1)}-c_{\phi q}^{(3)}$ is untestable with EWPO but only if it holds right at the $Z$ pole; RGE removes this one. as they are ubiquitous in models of new physics, arise easily at tree level in UV completions of the SMEFT and, with the exception of $\mathcal{O}_{ll}$ with first family leptons, none enters into EWPO at tree level [deBlas:2015aea]. The most unconstrained four-fermions involve third-family fermions only, in which we concentrate. Theoretically, one can motivate that new physics couples mostly to the third generation [Contino:2003ve, Panico:2015jxa, Davighi:2023iks, Chung:2023gcm]. We also ignore $c_{quqd}^{(1)}$, $c_{quqd}^{(8)}$, $c_{ledq}$, $c_{lequ}^{(1)}$ and $c_{lequ}^{(3)}$, as they are in general flavour-violating. We follow the notation of Ref. [Grzadkowski:2010es] for SMEFT WCs.
@@
We assume WCs to be defined at a scale $\Lambda=1$ TeV, and study their impact on EWPO at the scale $\sim m_{Z}$ using the leading-log dimension-six SMEFT RGEs, as well as upon integrating numerically these RGEs. To this latter goal, we rely on smelli which is further based on flavio [Straub:2018kue] and Wilson [Aebischer:2018bkb].
@@
To start with, the following WCs do not contribute to EWPO or their contribution is negligible: $c_{uu}$, $c_{dd}$, $c_{qu}^{(8)}$, $c_{qd}^{(8)}$, $c_{ud}^{(8)}$, $c_{le}$, $c_{ll}$, $c_{ed}$, $c_{ld}$. This explains why most of these WCs do not appear in Fig. 1 of Ref. [Maura:2024zxz]. Out of these, though, $c_{uu}$ is known to be well constrained by two-loop effects captured by RGE re-summation [Allwicher:2023aql, Allwicher:2023shc, Stefanek:2024kds, Maura:2024zxz]; see Fig. 1.
@@
where we have included the EWPOs that receive the largest contributions from these WCs. The null vector of the corresponding matrix, which describes the blind direction in this sub-space, is given by: $c_{ud}^{(1)}\sim c_{qd}^{(1)}$. It extends to the full set of EWPOs.
@@
$\Xi_{1}$ $(1,3)_{1}$ $\frac{|y_{\Xi_{1}}|^{2}}{M_{\Xi_{1}}^{2}}\mathcal{O}_{ll}$
$\varphi$ $(1,2)_{1/2}$ $\frac{|y^{e}_{\varphi}|^{2}}{2M_{\varphi}^{2}}\mathcal{O}_{le}-\frac{|y^{d}_{\varphi}|^{2}}{6M_{\varphi}^{2}}\left(\mathcal{O}^{(1)}_{qd}+6\mathcal{O}^{(8)}_{qd}\right)$
$\omega_{1}$ $(3,1)_{-1/3}$ $\frac{|y^{du}_{\omega_{1}}|^{2}}{3M_{\omega_{1}}^{2}}\left(\mathcal{O}^{(1)}_{ud}-3\mathcal{O}^{(8)}_{ud}\right)+\frac{|y^{eu}_{\omega_{1}}|^{2}}{2M_{\omega_{1}}^{2}}\mathcal{O}_{eu}$
$\frac{|y^{ql}_{\omega_{1}}|^{2}}{4M_{\omega_{1}}^{2}}\left(\mathcal{O}^{(1)}_{lq}-\mathcal{O}^{(3)}_{lq}\right)$
$\omega_{4}$ $(3,1)_{-4/3}$ $\frac{|y^{ed}_{\omega_{4}}|^{2}}{2M_{\omega_{4}}^{2}\mathcal{O}_{ed}}$
$\Pi_{1}$ $(3,2)_{1/6}$ $\frac{|y_{\Pi_{1}}|^{2}}{2M_{\Pi_{1}}^{2}}\mathcal{O}_{ld}$
$\Pi_{7}$ $(3,2)_{7/6}$ $\frac{|y^{eq}_{\Pi_{7}}|^{2}}{2M_{\Pi_{7}}^{2}}\mathcal{O}_{qe}-\frac{|y^{lu}_{\Pi_{7}}|^{2}}{2M_{\Pi_{7}}^{2}}\mathcal{O}_{lu}$
$\zeta$ $(3,3)_{-1/3}$ $\frac{|y^{ql}_{\zeta}|^{2}}{4M_{\zeta}^{2}}\left(3\mathcal{O}^{(1)}_{lq}+\mathcal{O}^{(3)}_{lq}\right)$
$\Omega_{1}$ $(6,1)_{1/3}$ $\frac{|y^{ud}_{\Omega_{1}}|^{2}}{6M_{\Omega_{1}}^{2}}\left(2\mathcal{O}^{(1)}_{ud}+3\mathcal{O}^{(8)}_{ud}\right)$
$\Omega_{2}$ $(6,1)_{-2/3}$ $\frac{|y_{\Omega_{2}}|^{2}}{2M_{\Omega_{2}}^{2}}\mathcal{O}_{dd}$
$\Phi$ $(8,2)_{1/2}$ $\frac{|y^{dq}_{\Phi}|^{2}}{18M_{\Phi}^{2}}\left(4\mathcal{O}^{(1)}_{qd}-3\mathcal{O}^{(8)}_{qd}\right)$
@@
$\mathcal{B}$ $(1,1)_{0}$ $\frac{|g^{d}_{\mathcal{B}}|^{2}}{2M_{\mathcal{B}}^{2}}\mathcal{O}_{dd}\,-\frac{|g^{l}_{\mathcal{B}}|^{2}}{2M_{\mathcal{B}}^{2}}\mathcal{O}_{ll}$
$\mathcal{B}_{1}$ $(1,1)_{1}$ $\frac{|g^{du}_{\mathcal{B}_{1}}|^{2}}{3M_{\mathcal{B}_{1}}^{2}}\left(\mathcal{O}^{(1)}_{ud}+6\mathcal{O}^{(8)}_{ud}\right)$
$\mathcal{W}$ $(1,3)_{0}$ $\frac{|g^{l}_{\mathcal{W}}|^{2}}{8M_{\mathcal{W}}^{2}}\mathcal{O}_{ll}$
$\mathcal{G}$ $(8,1)_{0}$ $\frac{|g^{d}_{\mathcal{G}}|^{2}}{4M_{\mathcal{G}}^{2}}\mathcal{O}_{dd}$
$\mathcal{G}_{1}$ $(8,1)_{1}$ $\frac{|g_{\mathcal{G}_{1}}|^{2}}{9M_{\mathcal{G}_{1}}^{2}}\left(4\mathcal{O}^{(1)}_{ud}-3\mathcal{O}^{(8)}_{ud}\right)$
$\mathcal{L}_{3}$ $(1,2)_{-3/2}$ $\frac{|g_{\mathcal{L}_{3}}|^{2}}{M_{\mathcal{L}_{3}}^{2}}\mathcal{O}_{le}$
$\mathcal{U}_{2}$ $(3,1)_{2/3}$ $\frac{|g^{ed}|^{2}}{M_{\mathcal{U}_{2}}^{2}}\mathcal{O}_{ed}-\frac{|g^{lq}_{\mathcal{U}_{2}}|^{2}}{2M_{\mathcal{U}_{2}}^{2}}\left(\mathcal{O}^{(1)}_{lq}+\mathcal{O}^{(3)}_{lq}\right)$
$\mathcal{U}_{5}$ $(3,1)_{5/3}$ $\frac{|g_{\mathcal{U}_{5}}|^{2}}{M_{\mathcal{U}_{5}}^{2}}\mathcal{O}_{eu}$
$\mathcal{Q}_{1}$ $(3,2)_{1/6}$ $\frac{2|g^{dq}_{\mathcal{Q}_{1}}|^{2}}{3M_{\mathcal{Q}_{1}}^{2}}\left(\mathcal{O}^{(1)}_{qd}-3\mathcal{O}^{(8)}_{qd}\right)+\frac{|g^{ul}_{\mathcal{Q}_{1}}|^{2}}{M_{\mathcal{Q}_{1}}^{2}}\mathcal{O}_{lu}$
$\mathcal{Q}_{5}$ $(3,2)_{-5/6}$ $\frac{|g^{dl}_{\mathcal{Q}_{5}}|^{2}}{M_{\mathcal{Q}_{5}}^{2}}\mathcal{O}_{ld}+\frac{|g^{eq}_{\mathcal{Q}_{5}}|^{2}}{M_{\mathcal{Q}_{5}}^{2}}\mathcal{O}_{qe}$
$\mathcal{X}$ $(3,3)_{2/3}$ $\frac{|g_{\mathcal{X}}|^{2}}{8M_{\mathcal{X}}^{2}}\left(3\mathcal{O}^{(1)}_{lq}-\mathcal{O}^{(3)}_{lq}\right)$
$\mathcal{Y}_{1}$ $(\bar{6},2)_{1/6}$ $\frac{|g_{\mathcal{Y}_{1}}|^{2}}{3M_{\mathcal{Y}_{5}}^{2}}\left(2\mathcal{O}^{(1)}_{qd}+3\mathcal{O}^{(8)}_{qd}\right)$
@@
$\{$ $\mathcal{O}_{dd},\mathcal{O}_{qu}^{(8)},\mathcal{O}_{qd}^{(8)},\mathcal{O}_{ud}^{(8)},\mathcal{O}_{le},\mathcal{O}_{ll},\mathcal{O}_{ed},\mathcal{O}_{ld},$